Vector has y-component Ay= +13.0m. A makes an
angle of 32.0* counterclockwise from the +y-axis. (a) What is the
x-component of A ? (b) What is the magnitude of A ?
1
Expert's answer
2019-10-11T10:52:34-0400
"A_x=-A_y\tan{32}=-13\tan{32}=-8.12\ m"

"A=\frac{A_y}{\cos{32}}=\frac{13}{\cos{32}}=15.3\ m"